40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 23 21:32:43 2013, Koji, Update, IOO, Found the cause of mysterious MC motion 
    Reply  Wed Jul 24 14:26:24 2013, rana, Update, IOO, Found the cause of mysterious MC motion 
       Reply  Wed Aug 21 19:26:19 2013, rana, Update, IOO, Found the cause of mysterious MC motion 
Message ID: 8913     Entry time: Tue Jul 23 21:32:43 2013     Reply to this: 8917
Author: Koji 
Type: Update 
Category: IOO 
Subject: Found the cause of mysterious MC motion 

Thesedays we were continuously annoyed by unELOGGED activities of the interferometer.

MC2 LOCKIN was left on and has continuously injected frequency noise and beam pointing modulation
during all of the comissioning / vent preparation.

C1:SUS-MC2_LOCKIN2_OSC_FREQ was 0.075
C1:SUS-MC2_LOCKIN2_OSC_CLKGAIN was 99

For more than a week ago we noticed that the curve of the MC WFS stripchart suddenly got THICKER.
MC WFS, arm transmission, beam pointing... everything was modulated.
It was not WFS instability, and it was not the cavity mirrors.

Today I made the investigation and finally tracked down the cause of this issue to be on MC2 suspension.
Then it was found that this LOCKIN was ON.

There is no direct record of this lockin in the frame files.
From the recorded channel "C1:IOO-WFS2-YAW_OUT16" (which is the trace on the StripTool chart on the wall)
It was turned on at July 10th, 2:00UTC (July 9th, 7PM PDT)

ELOG V3.1.3-